• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. Aniol

    Log entry time 01:51:53 on December 13,1999

    Entry number 30912

    keyword=kin #17, Help!

    At 00:30 we gave up trying to get kinematics #17 on the empty cell. At 22:30 the crew started to set the hrse dipole but was unable to do so . At 00:30 we decided to go to the ep settings on hydrogen. The problem with kinematics 17 is that the field in hrse is about 0.15T, but the probes only read down to 0.17T. We were unable to find an nmr lock. This problem was discussed with M. Kuss, who said that J. Lerose had previously offered to help establish these low fields.